# Multi-GPU Quantum Circuit Simulation and the Impact of Network Performance

### v1

@@ 2.6 Benchmark Systems and Methodology
The primary system used for the studies is “Genesis”, an in-house GB200 NVL72 rack. The system has 4 GB200 “Blackwell” GPUs and 2 NVIDIA Grace CPUs per node with 72 GPUs connected by all-to-all MNNVL generation 5 fabric. In addition to MNNVL, the system is configured with one ConnectX-7 (NDR InfiniBand) 400Gb port per GPU with a fully-balanced, non-blocking fat tree topology. For the previous generation systems, we use “Hopper” to refer to a single x86 server with a single 80GB HBM3 H100 GPU. “Ampere” refers to the generation before “Hopper” with eight 80GB A100 GPUs in a single x86 server connected by NVL. “Ampere-PCI” refers to a single x86 server with eight 80GB A100 GPUs with a PCI-express interconnect. As a baseline multi-node system using A100 GPUs, we use the Perlmutter system at the National Energy Research Scientific Computing Center (NERSC). Perlmutter has 4 A100 GPUs per node with one Slingshot 11 “Cassini” NIC per GPU using a 3-hop dragonfly topology for the internode network [perlmutter]. Over 7000 GPUs are available on the entire system. Further details of the systems are available in Table 3.
@@ 3 Results
Single GPU generational speedups for the 33-qubit QPE and HamLib benchmarks are shown in Figure 1 with approximately 1.8X speedup going from the NVIDIA Ampere to Hopper generations and 2.2-2.4X from Hopper to Blackwell. Multi-GPU weak scaling performance is shown in Figure 2, starting with 33 qubits and increasing this number by one with each doubling of GPU counts. This represents the largest simulations fitting into GPU memory on Perlmutter. Due to the higher memory capacity of the GB200 GPUs, we could simulate between 34 and 40 qubits on the Genesis system with better parallel efficiency. However, this would obfuscate the comparison to previous generations. The ideal weak scaling performance is based on simulation time for a single GB200 GPU, normalized to account for the increasing gate count for larger circuits with more qubits.
@@
For the 33-qubit HamLib benchmark, we observe lower network sensitivity (Figure 4). Performance is monotonically improving in all cases, however, and MNNVL still outperforms InfiniBand by 1.5-3X on Genesis. Although bisection-BW is the primary concern here, we still observe a significant performance decrease when disabling RDMA between the GPUs and InfiniBand NICs (Genesis-IB-RDMA vs Genesis IB). In the strong-scaling cases, the performance hit ranges from 13% to 59%; with weak-scaling, it can be as high as 68%. In comparing MNNVL algorithms in cuQuantum, implementations with the low-level VMM API do significantly outperform CUDA-aware MPI. Here, the impact increases at higher GPU counts. For HamLib, this ranged from 1.06X to 1.25X between 2 and 16 nodes. For QPE with strong scaling, the range was 1.11-1.41X, and for weak scaling, the range was 1.11-1.61X. We expect a larger impact from algorithmic differences here compared to the benefits of zero-copy without buffering. For single-node benchmarks, the NVL interconnect significantly outperformed PCI-express with up to 2.5X performance for the HamLib benchmark and 3.5X for QPE.
@@
In comparing Genesis to Perlmutter, intranode speedups with 1-4 GPUs were balanced and did not significantly deviate from those achieved with a single GPU. Disproportionate advances in the internode interconnect with MNNVL yield larger speedups as the node count increases. Strong-scaling performance was between 9.7X and 15X higher for multiple nodes on Genesis for QPE and 6X and 8.1X higher for HamLib. Weak scaling performance was 13X at 36 qubits and 16.2X-16.8X at higher qubit counts. Figure 5 illustrates the speedups compared to Perlmutter, with the best simulation times for each configuration tested here.
@@ 4 Conclusion
While our focus here has been on HW and APIs for communication, algorithmic advances will also be critical. In this regard, sophisticated optimizations for the NVIDIA Grace-Hopper architecture were employed in the Jülich universal quantum computer simulator, demonstrating a capability for 50-qubit simulations when using the entire supercomputer [deraedt2025universalquantumsimulation50]. Of course, there are many other examples with ongoing work in-house and elsewhere [jiao2023communication, rezaei2025low, teranishi2025lazy, gangapuram2024benchmarking, stenzel2024qandle, westrick2024grafeyn, adamski2023energy]. We expect continued improvements, both in algorithms and HW, that will prove vital in advancing the field of quantum computing.
@@ 5 Acknowledgments
We thank Takuma Yamaguchi and Akshay Venkatesh for their expert reviews of this work.

### v2

@@ 2.6 Benchmark Systems and Methodology
The primary system used for the studies is “Genesis”, an in-house GB200 NVL72 rack. The system has 4 GB200 “Blackwell” GPUs and 2 NVIDIA Grace CPUs per node with 72 GPUs connected by all-to-all MNNVL generation 5 fabric. In addition to MNNVL, the system is configured with one ConnectX-7 (NDR InfiniBand) 400Gb port per GPU with a fully-balanced, non-blocking fat tree topology. For the previous generation systems, we use “Hopper” to refer to a single x86 server with a single 80GB HBM3 H100 GPU. “Ampere” refers to the generation before “Hopper” with eight 80GB A100 GPUs in a single x86 server connected by NVL. “Ampere-PCI” refers to a single x86 server with eight 80GB A100 GPUs with a PCI-express interconnect. As a baseline multi-node system using A100 GPUs, we use the Perlmutter system at the National Energy Research Scientific Computing Center (NERSC). Perlmutter has 4 A100 GPUs per node with one Slingshot 11 “Cassini” NIC per GPU using a 3-hop dragonfly topology for the internode network [38]. Over 7000 GPUs are available on the entire system. Further details of the systems are available in Table 3.
@@
2.7 Profiling
Currently, the MPI implementation in cuQuantum uses dynamic library loading with direct calls to implementation-specific MPI functions. This is intended to prevent a requirement for recompilation with specific MPI configurations on a given system. However, because it bypasses direct use of the MPI API, this approach may prevent accurate use of profiling tools that intercept MPI functions at a higher level. Therefore, we use the perf tool for profiling with the --sort dso option for reporting. This allows identification of the total percentage of simulation time spent in MPI. As the current implementation does not overlap communication and computation for state-vector simulation, the approach is viable without additional steps to separate exposed communication time. The --call-graph dwarf option for profiling is used to capture non-self time from calls into other libraries. To force all communications through MPI, the low-level APIs are disabled (accomplished by setting the environment variable CUDAQ_GPU_FABRIC=1). To avoid profiling time in warmup runs, the --delay option is used with the delay time determined from the end-to-end time for simulation of a single run. The number of circuits simulated was increased to ensure at least a 60-second run time for profiling. We advise that care should be taken to evaluate any profiling overhead in the simulation time. We were unable to apply this approach to the HamLib benchmark due to excessive profiling overhead (more than 2X simulation time) associated with routines handling the JIT-compiled CUDA-Q kernels.
@@ 3 Results
Single GPU generational speedups for the 33-qubit QPE, HamLib, and RCS benchmarks are shown in Figure 1 with approximately 1.8X speedup going from the NVIDIA Ampere to Hopper generations and 2.2-2.4X from Hopper to Blackwell. Multi-GPU weak scaling performance for the QPE benchmark is shown in Figure 2, starting with 33 qubits and increasing this number by one with each doubling of GPU counts. This represents the largest simulations fitting into GPU memory on Perlmutter. Due to the higher memory capacity of the GB200 GPUs, we could simulate between 34 and 40 qubits on the Genesis system with better parallel efficiency. However, this would obfuscate the comparison to previous generations. The ideal weak scaling performance is based on simulation time for a single GB200 GPU, normalized to account for the increasing gate count for larger circuits with more qubits.
@@
For the 33-qubit HamLib benchmark, we observe lower network sensitivity (Figure 4). For a given qubit count on a specific system, lower network sensitivity for a simulation might occur for two reasons. The first is increased time in GPU kernels for gate application, where specialized kernels apply different types of gates at different fusion levels, resulting in different execution times per gate. The second is a circuit whose structure allows fewer elements to be exchanged per gate between GPUs during simulation. In this HamLib benchmark, the lower sensitivity is due to the latter. The TFIM Hamiltonian, given by equation 1, describes interacting particles as a linear chain with nearest-neighbor interactions. This is translated into a circuit with a ladder-like structure, allowing more gate applications to be performed without exchanging elements between MPI tasks. This results in less time spent in inter-GPU communication and thus less potential improvement from network improvements. Performance is monotonically improving in all cases, however, and MNNVL still outperforms InfiniBand by 1.5-3X on Genesis.
In contrast to the QPE and HamLib benchmarks, the random circuits exhibit irregular structure and connectivity by construction. This can be a useful test, as it is unexpected that optimizations from general distributed state vector algorithms would allow an accurate solution. When benchmarking the random samples, we observe very similar scaling behavior and parallel efficiency to those of the QPE results. MNNVL speedups over InfiniBand ranged from 2-3.71X for weak scaling and from 2.13-3.24X for the 33-qubit circuits.
@@
Although bisection-BW is the primary concern here, we still observe a significant performance decrease when disabling RDMA between the GPUs and InfiniBand NICs (Genesis-IB-RDMA vs Genesis IB). In the strong-scaling cases, the performance hit ranges from 13% to 59%; with weak-scaling, it can be as high as 68%. In comparing MNNVL algorithms in cuQuantum, implementations with the low-level VMM API do significantly outperform CUDA-aware MPI. Here, the impact increases at higher GPU counts. For HamLib, this ranged from 1.06X to 1.25X between 2 and 16 nodes. For QPE with strong scaling, the range was 1.11-1.41X, and for weak scaling, the range was 1.11-1.61X. We expect a larger impact from algorithmic differences here compared to the benefits of zero-copy without buffering. For single-node benchmarks, the NVL interconnect significantly outperformed PCI-express with up to 2.5X performance for the HamLib benchmark and 3.5X for QPE.
In terms of expected speedups from the network performance improvements, we can estimate an idealized speedup by profiling to determine the fraction of simulation time () spent in MPI. In this case, the ideal speedup is given by where is the inverse ratio of the bisection bandwidths. This estimate can be complicated by hierarchical communication when using the low-level API for intranode NVL communication. However, for QPE, we observe similar performance on the Genesis system with this disabled, so all inter-GPU communication is handled through the MPI library. This allows a simple method for profiling the InfiniBand performance. These results are shown in Table 4 for the QPE benchmark on the Genesis system. In this case, we compare measured MNNVL-CUDA speedups with ideal speedups from profiling in the IB-RDMA configuration with hierarchical communications disabled (i.e., disabling the use of the low-level API by setting CUDAQ_GPU_FABRIC). For this system, , and the measured speedups are from comparison to performance using the InfiniBand internode network with only MPI (results not plotted) to account for any small performance changes from previously shown results.
| Nodes | 2 | 4 | 8 | 16 |
| -Qubit %MPI | 68% | 74% | 76% | 78% |
| Ideal Speedup | 2.9 | 3.6 | 3.8 | 4.1 |
| Measured | 2.6 | 3.4 | 3.6 | 4.0 |
| 33-Qubit %MPI | 65% | 72% | 70% | 64% |
| Ideal Speedup | 2.7 | 3.3 | 3.1 | 2.6 |
| Measured | 2.5 | 3.7 | 2.8 | 2.8 |
Deviations from the ideal speedups (in either direction) might occur from non-uniform percentages of peak network BW, factors outside of network BW, imperfect non-network load balance, variations in hardware, and volatile state of the nodes, latencies within the communication library code, communication buffering, and the approximate nature of the profiling method. However, we observe good agreement with expectations for the algorithm as shown in Table 4. In support of the methodology, the 33-qubit and 39-qubit times with Infiniband are 0.843s and 6.965s giving an approximate parallelization cost of , or 83% of the simulation time. While Table 4 indicates 78% of time in MPI, we expect additional parallelization overhead outside of MPI. For the InfiniBand configuration, all cases are MPI-bound with over 50% of time spent in the library, and the speedups match expectations based on the bisection BWs. With a 36X performance improvement in BW, performance can become sensitive to implementation details in inter-GPU communication. As observed with the lower performance of the CUDA-aware MPI implementation shown in the earlier results, the low-level CUDA implementation is currently required to realize the full potential speedups. The lower percentage of time spent in MPI with strong scaling for the 33-qubit benchmark indicates lower parallel performance due to increased time spent in serial code paths and parallelization overheads that are not directly related to network performance.
@@
In comparing Genesis to Perlmutter, intranode speedups with 1-4 GPUs were balanced and did not significantly deviate from those achieved with a single GPU. Disproportionate advances in the internode interconnect with MNNVL yield larger speedups as the node count increases. Strong-scaling performance was between 9.7X and 15X higher for multiple nodes on Genesis for QPE and 6X and 8.1X higher for HamLib. Weak scaling performance was 13X at 36 qubits and 16.2X-16.8X at higher qubit counts. These results are somewhat higher than expectations from peak BWs, MPI profiling on Genesis (78% at 39-qubits), and the single-GPU generational speedup (4.51X) that would give an expected approximate speedup of . This indicates a lower percentage of peak interconnect BW achieved for simulations on the system, supported by a measured 77.5% of time in MPI on Perlmutter for 39 qubits. As with the QPE benchmark, for RCS we observed higher speedups with 64 GPUs at larger qubit counts with 14.1X speedup at 39 qubits. The 33-qubit speedup is still very significant at 11.7X. Figure 5 illustrates the speedups compared to Perlmutter, with the best simulation times for each configuration tested here.
@@ 4 Conclusion
While our focus here has been on HW and APIs for communication, algorithmic advances will also be critical. In this regard, sophisticated optimizations for the NVIDIA Grace-Hopper architecture were employed in the Jülich universal quantum computer simulator, demonstrating a capability for 50-qubit simulations when using the entire supercomputer [50]. Of course, there are many other examples with ongoing work in-house and elsewhere [25, 51, 59, 20, 56, 65, 1]. We expect continued improvements, both in algorithms and HW, that will prove vital in advancing the field of quantum computing.
@@ 5 Acknowledgments
We thank Takuma Yamaguchi and Akshay Venkatesh for their expert reviews of this work. Likewise, we thank the reviewers for their time and effort to provide feedback to improve the quality of the paper.